Almost all companies utilize some type of​ year-end performance review for their employees. Human Resources​ (HR) at a​ university's Health Science Center provides guidelines for supervisors rating their subordinates. For​ example, raters are advised to examine their ratings for a tendency to be either too lenient or too harsh. According to​ HR, "if you have this​ tendency, consider using a normal distributionlong dash—​10% of employees​ (rated) exemplary,​ 20% distinguished,​ 40% competent,​ 20% marginal, and​ 10% unacceptable." Suppose you are rating an​ employee's performance on a scale of 1​(lowest) to 100​ (highest). Also, assume the ratings follow a normal distribution with a mean of 51 and a standard deviation of 17.Complete parts a and b.

b. What is the lowest rating you should give to a​ "competent" employee if you follow the​ university's guidelines?

____________​(Round to two decimal places as​ needed.)
Answer #1

here as we can see that lowest rating for a competent" employee is at 30th percentile for which

z score =-0.52

therefore corresponding score=mean+z*std deviation =51-0.52*17=42.16